The National Football League (NFL) records a variety of performance data for individuals and teams. To investigate the importance of passing on the percentage of games won by a team, the following data show the conference (Conf), average number of passing yards per attempt (Yds/Att), the number of interceptions thrown per attempt (Int/Att), and the percentage of games won (win96) for a random sample of 16 NFL teams for a season Conf NFC NFC NFC AFC NFC NFC AFC AFC AFC NFC AFC NFC AFC NFC AFC NFC Yds/Att 6.4 7.1 7.3 6.3 7.3 8.9 7.5 5.8 4.7 5.9 8.4 8.1 7.4 6.7 6.7 6.3 Int/Att 0.042 0.023 0.034 0.026 0.025 0.013 0.019 0.026 0.034 0.032 0.020 0.023 0.045 0.011 0.024 0.041 Team Arizona Cardinals Atlanta Falcons Carolina Panthers Cincinnati Bengal:s Detroit Lions Green Bay Packers Houstan Texans Indianapolis Colts Jacksonville Jaguars Minnesota Vikings New England Patriots New Orleans Saints Oakland Raiders San Francisco 49ers Tennessee Titans Washington Redskins Win % 49.9 62.2 37.6 56.3 62.2 94.0 62.6 12.2 31.4 18.9 81.2 81.4 50.1 81.1 56.1 31.0
a. Develop the estimated regression equation that could be used to predict the percentage of games won given the average number of passing yards per attempt to 1 decimal). Enter negative value as negative number Win% b. Develop the estimated regression equation that could be used to predict the percentage of games won glven the number of interceptions thrown per attempt to 1 decimal). Enter negative value as negative number Win% Yds/Att Int/Att c.Develop the estimated regression equation that could be used to predict the percentage of games won given the average number of passing yards per attempt and the number of interceptions thrown per attempt (to 1 decimal). Enter negative value as negative number Win% d. The average number of passing yards per attempt for the Kansas City Chiefs was 6.2 and the number of interceptions thrown per attempt was 0.036. Use the estimated regression equation developed in part (c) to predict the percentage of games won by the Kansas City Chiefs. (Note: For a season the Kansas City Chiefs' record was 7 wins and 9 losses.) Compare your prediction to the actual percentage of games won by the Kansas City Chiefs (to whole number) Predicted percentage .Yds/Att+ Int/Att Actual percentage Select your answer
Answer #1

a)

The regression equation is
Win% = - 67.5 + 17.6 Yds/Att

b)


The regression equation is
Win% = 81.1 - 1053.4 Int/Att

c)

The regression equation is
Win% = - 43.5 + 16.0 Yds/Att - 519.3 Int/Att

d)

Predicted

Win% = - 43.5 + 16.0*6.2 - 519.3*0.036=37

Actual percentage of win=(7/(7+9))*100%=44%

Predicted percentage is less than actual percentage.